# Tightening the entropic uncertainty relations with quantum memory in a multipartite scenario

### v1

@@ II Tripartite QMA-EUR and generalized QMA-EURs for multiple measurements in multipartite systems
Figure 1: The horizontal axis represents the parameter $a$, while the vertical axis represents the difference between our result and the previous results. This difference (red curve) remains above the horizontal axis, indicating that our results are consistently better than the previous ones.
From Theorem 2 and 3 we deduce the following optimal lower bound,
$\sum_{t=1}^{n}\sum_{M_{i}\in\mathbf{S}_{t}}S(M_{i}|B_{t})\geq$ $\frac{1}{m-1}\sum_{i<j}Q^{M_{i,j}}+\sum_{t=1}^{n}\frac{m_{t}(m_{t}-1)}{2(m-1)}S(A|B_{t})+\max\{0,\delta_{mn},\delta_{mn}^{\prime\prime}\}.$ (21)
@@
$\sum_{i=1}^{m}S\left(M_{i}\mid B\right)\geq$ $\frac{1}{m-1}\sum_{i<j}Q^{M_{i,j}}+\frac{m}{2}S(A|B)+\max\left\{0,\delta_{m1},\delta_{m1}^{{}^{\prime\prime}}\right\},$
@@
$\delta_{m1}^{{}^{\prime\prime}}=-\frac{1}{m}\omega\beta+(m-1)S(A)-\frac{m}{2}S(A)+\frac{m}{2}\mathcal{I}(A:B)-\frac{1}{m-1}\sum_{i<j}Q^{M_{i,j}}-\sum_{i}^{m}\mathcal{I}(M_{i}:B).$
@@
For the case of $m=n$, $m$ memories with the cardinality one of $\mathbf{S}_{t}$ ($m_{t}=1$), the inequality (21) reduces to
@@
$\delta_{mm}^{{}^{\prime\prime}}=-\frac{1}{m}\omega\beta+(m-1)S(A)-\sum_{i}^{m}\mathcal{I}(M_{i}:B_{i})-\frac{1}{m-1}\sum_{i<j}Q^{M_{i,j}}.$
@@ III Multipartite QMA-EURs for POVMs
where $q^{X_{1,2}}(\rho_{A})=\mathrm{max}\{q(\rho_{A},X_{1},X_{2}),q(\rho_{A},X_{2},X_{1})\}$ with $q(\rho_{A},X_{1},X_{2})=-\sum\limits_{j}p^{X_{1}}_{j}\log_{2}\\ h_{j}(X_{1},X_{2})$ and $q(\rho_{A},X_{2},X_{1})=-\sum\limits_{k}p^{X_{2}}_{k}\log_{2}h_{k}(X_{2},X_{1})$. Here, $p^{X_{1}}_{j}=\mathrm{tr}(X^{1}_{j}\rho_{A})$, $p^{X_{2}}_{k}=\mathrm{tr}(X^{2}_{k}\rho_{A})$ and $h_{j}(X_{1},X_{2})=\|\sum\limits_{k}X^{2}_{k}X^{1}_{j}X^{2}_{k}\|_{\infty}$, $h_{k}(X_{2},X_{1})=\|\sum\limits_{j}X^{1}_{j}X^{2}_{k}X^{1}_{j}\|_{\infty}$, where $\|\cdot\|_{\infty}$ denotes the infinity norm, $f^{X_{1,2}}=\mathrm{min}\{S(A|BX_{1}),S(A|BX_{2})\}$, $S(A|BX_{1})$ and $S(A|BX_{2})$ are the conditional von Neumann entropies of ${\rho}_{X_{1}AB}$ and ${\rho}_{X_{2}AB}$, respectively Coles and Piani (2014). Notably, for orthonormal bases (rank-one projective POVMs), $f^{X_{1,2}}=0$. In the absence of quantum memory $B$, (22) reduces to
@@
$\sum_{t=1}^{n}\sum_{X_{i}\in\mathbf{S}_{t}}S(X_{i}|B_{t})\geq$ $\frac{1}{m-1}\sum_{t=1}^{n}\frac{m_{t}(m_{t}-1)}{2}S(A|B_{t})+\frac{1}{m-1}\sum_{i<j}q^{X_{i,j}}(\rho_{A})$ (25)
$-\frac{1}{m-1}\sum_{t=1}^{n}\mathop{\sum}_{\begin{subarray}{c}X_{i^{\prime}},X_{j^{\prime}}\in S_{t}\\ i^{\prime}<j^{\prime}\end{subarray}}f^{X_{i^{\prime},j^{\prime}}}+\max\{0,\kappa_{mn}\},$
@@
$\kappa_{mn}=$ $\frac{m(m-1)-\sum\limits_{t=1}^{n}m_{t}(m_{t}-1)}{2(m-1)}S(A)+\sum_{t=1}^{n}\frac{m_{t}(m_{t}-1)}{2(m-1)}\mathcal{I}(A:B_{t})-\sum_{t=1}^{n}\sum_{X_{i}\in\mathbf{S}_{t}}\mathcal{I}(X_{i}:B_{t})$
@@
$\sum_{t=1}^{n}\sum_{X_{i}\in\mathbf{S}_{t}}S(X_{i}|B_{t})\geq$ $\frac{1}{m-1}\sum_{i<j}q^{X_{i,j}}(\rho_{A})+\frac{m}{2}S(A)-\sum_{t=1}^{n}\sum_{X_{i}\in\mathbf{S}_{t}}\mathcal{I}(X_{i}:B_{t})-\frac{1}{m-1}\sum_{i<j}^{m}\tilde{f}^{X_{i,j}}.$ (26)
@@
$S(X_{i}|B_{t})+S(X_{j}|B_{s})\geq\left\{\begin{aligned} &q^{X_{i,j}}(\rho_{A}),\quad t\neq s\\ &q^{X_{i,j}}(\rho_{A})+S(A|B_{t})-f^{X_{i,j}},\quad t=s.\end{aligned}\right.$
@@
$\sum_{t=1}^{n}\sum_{X_{i}\in\mathbf{S}_{t}}S(X_{i}|B_{t})\geq$ $\frac{1}{m-1}\sum_{t=1}^{n}\frac{m_{t}(m_{t}-1)}{2}S(A|B_{t})+\frac{1}{m-1}\sum_{i<j}q^{X_{i,j}}(\rho_{A})$ (27)
$-\frac{1}{m-1}\sum_{t=1}^{n}\mathop{\sum}_{\begin{subarray}{c}X_{i^{\prime}},X_{j^{\prime}}\in S_{t}\\ i^{\prime}<j^{\prime}\end{subarray}}f^{X_{i^{\prime},j^{\prime}}},$
@@
$\sum_{i=1}^{m}S\left(X_{i}\mid B\right)\geq$ $\frac{1}{m-1}\sum_{i<j}q^{X_{i,j}}(\rho_{A})+\frac{m}{2}S(A|B)-\frac{1}{m-1}\sum_{i<j}f^{X_{i,j}}+\max\left\{0,\kappa_{m1}\right\},$
where $\kappa_{m1}=\displaystyle\frac{m}{2}\mathcal{I}(A:B)-\sum_{i=1}^{m}\mathcal{I}\left(X_{i}:B\right)+\frac{1}{m-1}\sum_{i<j}^{m}\left[f^{X_{i,j}}-\tilde{f}^{X_{i,j}}\right]$.
If the cardinality of $\mathbf{S}_{t}$ is set to be $1$ ($m_{t}=1$, $n=m$), Theorem 4 further simplifies to
@@ IV.1 Uncertainty relation for quantum coherence
Coherence is a fundamental issue in quantum physics and a crucial resource for quantum information processing. In 2014 Baumgratz et al. Baumgratz et al. (2014) introduced a framework for quantifying coherence. For a fixed orthonormal basis $\{|i\rangle\}^{d}_{i=1}$ of the $d$-dimensional Hilbert space $\mathcal{H}$, density matrices that are diagonal in this basis are termed incoherent. We denote the set of incoherent quantum states as =$\mathcal{I}=\{\sigma|\sigma=\sum\limits_{i=1}^{d}\lambda_{i}|i\rangle\langle i|\}$.
For a bipartite system $AB$, an $A$-incoherent state (also known as an incoherent-quantum state Chitambar et al. (2016)) was proposed in Ma et al. (2016) with respect to the basis (${|i\rangle_{A}}$) of subsystem $A$ This is also referred to as unilateral coherence Ma et al. (2017); Dolatkhah et al. (2019). We denote the set of $A$-incoherent states as $\mathcal{I}_{B|A}$,
@@
$C^{B|A}_{r}(\rho_{AB})=\mathop{\mathrm{min}}\limits_{\sigma_{AB}\in\mathcal{I}_{B|A}}S(\rho_{AB}\|\sigma_{AB})=S(\rho_{AB}\|\Delta_{A}(\rho_{AB}))=S(\Delta_{A}(\rho_{AB}))-S(\rho_{AB}),$
@@
$\sum_{t=1}^{n}\sum_{M_{i}\in\mathbf{S}_{t}}C^{M_{i}}_{r}(\rho_{AB_{t}})\geq\frac{1}{m-1}\sum_{i<j}Q^{M_{i,j}}+\sum_{t=1}^{n}\frac{m_{t}(m_{t}-2m+1)}{2(m-1)}S(A|B_{t})+\max\{0,\delta_{mn},\delta_{mn}^{{}^{\prime\prime}}\}.$ (30)
Example 2 Consider the two-qubit mixed state $\rho_{AB}=p|\sigma\rangle\langle\sigma|+\frac{1-p}{4}\mathrm{I_{AB}}$, where $0\leq p\leq 1$, $|\sigma\rangle=\cos{\alpha}|0_{A}0_{B}\rangle+\sin{\alpha}|1_{A}1_{B}\rangle$ with $\alpha\in[0,2\pi)$. Let the three complementary observables measured on subsystem $A$ be the Pauli matrices, $M_{1}=\sigma_{x}$, $M_{2}=\sigma_{y}$ and $M_{3}=\sigma_{z}$. In this scenario ($m_{t}=m=3$, $n=1$), our inequality (30) gives rise to
$C^{M_{1}}_{r}(\rho_{AB})+C^{M_{2}}_{r}(\rho_{AB})+C^{M_{3}}_{r}(\rho_{AB})$ $\geq\frac{3}{2}-\frac{3}{2}S(A|B)+\max\{0,\delta_{31},\delta_{31}^{{}^{\prime\prime}}\},$ (31)
@@
$\delta_{31}^{{}^{\prime\prime}}=-\frac{1}{3}\omega\beta+\frac{1}{2}S(A)+\frac{3}{2}\mathcal{I}(A:B)-\frac{1}{2}\sum_{i<j}Q^{M_{i,j}}-\sum_{i}^{3}\mathcal{I}(M_{i}:B).$
In Fig. 2, the lower bound of the coherence-based uncertainty relation (31) is illustrated for different parameters $p$ and $\alpha$.
Figure 2: The black (sold) curve represents the value of the sum $C^{M_{1}}_{r}(\rho_{AB})+C^{M_{2}}_{r}(\rho_{AB})+C^{M_{3}}_{r}(\rho_{AB})$. The green (dashed) curves represent the lower bounds (Lb) in inequality (31). (a) $p=\frac{1}{3}$; (b) $\alpha=\frac{\pi}{8}$.

### v2

@@ II Tripartite QMA-EUR and generalized QMA-EURs for multiple measurements in multipartite systems
Figure 2: The horizontal axis represents the parameter $a$, while the vertical axis represents the difference (diff) between our result and the previous result. This difference (red solid curve) remains above the horizontal axis, indicating that our result is consistently better than the previous one.
It can be seen from the Figure $1$ and Figure $2$, $LB2>lb1$, $LB2>lb2$, $lb2>LB1$, $LB2>LB1$, that is, $LB2>lb2>LB1>lb1$. It means that our lower bound $LB2$ is tighter than the previous results $lb1$ and $lb2$ and $LB2>LB1$.
The above Examples $1$ and $2$ demonstrate that our lower bound, namely, the maximum of $LB1$ and $LB2$, is tighter than the previous lower bounds $lb1$ and $lb2$. They also show that our lower bounds $LB1$ and $LB2$ are complementary in certain cases.
In order to better illustrate our results, we consider the more general case, the random three-qutrit states, $\rho_{ABC}=\sum_{s=1}^{27}\lambda_{s}|\varphi_{s}\rangle\langle\varphi_{s}|$, where $\lambda_{s}$ and $|\varphi_{s}\rangle$ denote the $s$th eigenvalue and eigenvector of $\rho_{ABC}$, respectively. In addition, the eigenvalue $\lambda_{s}$ corresponds to the probability of the system being in the state $|\varphi_{s}\rangle$. And the normalized eigenvector $|\varphi_{s}\rangle$ can constitute an arbitrary unitary operation $\mathcal{V}=\{|\varphi_{1}\rangle,|\varphi_{2}\rangle,\cdots,|\varphi_{27}\rangle\}$. Therefore, we can obtain an arbitrary three-qutrit state by using an arbitrary set of probabilities $\lambda_{s}$ and an arbitrary unitary operator $\mathcal{V}$. Here we adopt an effective approach to generate random states. Firstly, we define a random function $\xi(\theta_{1},\theta_{2})$, which can generate a independent random real number in a closed interval $\theta_{1},\theta_{2}]$. The random number $p_{s}$ is then generated:
$p_{1}=\xi(0,1),~p_{s+1}=\xi(0,1)p_{s},$
where $s\in\{1,2,\cdots,26\}$. We obtain a set of probabilities in descent order:
$\lambda_{s}=\frac{p_{s}}{\sum_{s=1}^{27}p_{s}},~s=1,\dots,27.$
Utilizing the random function $\xi(-1,1)$ within a closed interval [1,1][-1,1], we randomly generate a twenty-seven-order real matrix $\mathcal{R}$. Based on the real matrix $\mathcal{R}$, a random Hermitian matrix can be obtained as $\mathcal{M}=D+(U^{T}+U)+\mathrm{i}(L^{T}+L)$, where $D$, $U$ and $L$, respectively, denote the diagonal, strictly upper and lower triangular parts of the real matrix $\mathcal{R}$, $U^{T}$ is the transpose of $U$. We obtain twenty-seven normalized eigenvectors $|\varphi_{s}\rangle$ of the Hermitian matrix $\mathcal{M}$ and a random unitary operator $\mathcal{V}$. Therefore, the random three-qutrit states are constructed.
As an illustration, we use 50 sets of three-qutrit states and three-dimensional random measurements to depict the corresponding uncertainty and our bounds in Figure 3. It is seen from Figure 3 that $LB1>lb1$ and $LB2>lb1$, while $LB1$ and $lb2$ have an uncertain relationship, and $LB2>lb2$. This shows that the maximum of our lower bounds ($LB1$ and $LB2$) outperforms the previous bounds. Additionally, the figure illustrates that $LB1$ and $LB2$ are complementary in certain cases.
Figure 3: The horizontal axis represents the number of random measurements and random states, while the vertical axis represents the difference (diff) between our results and the previous results. The magenta triangle, black star, blue circle, green diamond and the red square represent the values of LB2lb1LB2-lb1,LB2lb2LB2-lb2,LB1lb1LB1-lb1 and LB1LB2LB1-LB2, respectively.
Therefore, we deduce the following optimal lower bound,
$\sum_{t=1}^{n}\sum_{M_{i}\in\mathbf{S}_{t}}S(M_{i}|B_{t})\geq$ $\frac{1}{m-1}\sum_{i<j}Q^{M_{i,j}}$ (21)
$+$ $\sum_{t=1}^{n}\frac{m_{t}(m_{t}-1)}{2(m-1)}S(A|B_{t})$
$+$ $\max\{0,\delta_{mn},\delta_{mn}^{\prime\prime}\}.$
@@
$\sum_{i=1}^{m}S\left(M_{i}\mid B\right)\geq$ $\frac{1}{m-1}\sum_{i<j}Q^{M_{i,j}}+\frac{m}{2}S(A|B)$
$+$ $\max\left\{0,\delta_{m1},\delta_{m1}^{{}^{\prime\prime}}\right\},$
@@
$\delta_{m1}^{{}^{\prime\prime}}=$ $-\frac{1}{m}\omega\beta+(m-1)S(A)-\frac{m}{2}S(A)+\frac{m}{2}\mathcal{I}(A:B)$
$-\frac{1}{m-1}\sum_{i<j}Q^{M_{i,j}}-\sum_{i}^{m}\mathcal{I}(M_{i}:B).$
@@
For the case of $m=n$, $m$ memories with the cardinality of $\mathbf{S}_{t}$ one ($m_{t}=1$), the inequality (21) reduces to
@@
$\delta_{mm}^{{}^{\prime\prime}}=$ $-\frac{1}{m}\omega\beta+(m-1)S(A)-\sum_{i}^{m}\mathcal{I}(M_{i}:B_{i})$
$-\frac{1}{m-1}\sum_{i<j}Q^{M_{i,j}}.$
@@ III Multipartite QMA-EURs for POVMs
where $q^{X_{1,2}}(\rho_{A})=\mathrm{max}\{q(\rho_{A},X_{1},X_{2}),q(\rho_{A},X_{2},X_{1})\}$ with $q(\rho_{A},X_{1},X_{2})=-\sum\limits_{j}p^{X_{1}}_{j}\log_{2}h_{j}(X_{1},X_{2})$ and $q(\rho_{A},X_{2},X_{1})=-\sum\limits_{k}p^{X_{2}}_{k}\log_{2}h_{k}(X_{2},X_{1})$. Here, $p^{X_{1}}_{j}=\mathrm{tr}(X^{1}_{j}\rho_{A})$, $p^{X_{2}}_{k}=\mathrm{tr}(X^{2}_{k}\rho_{A})$ and $h_{j}(X_{1},X_{2})=\|\sum\limits_{k}X^{2}_{k}X^{1}_{j}X^{2}_{k}\|_{\infty}$, $h_{k}(X_{2},X_{1})=\|\sum\limits_{j}X^{1}_{j}X^{2}_{k}X^{1}_{j}\|_{\infty}$, where $\|\cdot\|_{\infty}$ denotes the infinity norm, $f^{X_{1,2}}=\mathrm{min}\{S(A|BX_{1}),S(A|BX_{2})\}$, $S(A|BX_{1})$ and $S(A|BX_{2})$ are the conditional von Neumann entropies of ${\rho}_{X_{1}AB}$ and ${\rho}_{X_{2}AB}$, respectively Coles and Piani (2014). Notably, for orthonormal bases (rank-one projective POVMs), $f^{X_{1,2}}=0$. In the absence of quantum memory $B$, (22) reduces to
@@
$\sum_{t=1}^{n}\sum_{X_{i}\in\mathbf{S}_{t}}S(X_{i}|B_{t})$ (25)
$\geq$ $\frac{1}{m-1}\sum_{t=1}^{n}\frac{m_{t}(m_{t}-1)}{2}S(A|B_{t})$
$+$ $\frac{1}{m-1}\sum_{i<j}q^{X_{i,j}}(\rho_{A})-\frac{1}{m-1}\sum_{t=1}^{n}\mathop{\sum}_{\begin{subarray}{c}X_{i^{\prime}},X_{j^{\prime}}\in S_{t}\\ i^{\prime}<j^{\prime}\end{subarray}}f^{X_{i^{\prime},j^{\prime}}}$
$+$ $\max\{0,\kappa_{mn}\},$
@@
$\kappa_{mn}=$ $\frac{m(m-1)-\sum\limits_{t=1}^{n}m_{t}(m_{t}-1)}{2(m-1)}S(A)$
$+$ $\sum_{t=1}^{n}\frac{m_{t}(m_{t}-1)}{2(m-1)}\mathcal{I}(A:B_{t})-\sum_{t=1}^{n}\sum_{X_{i}\in\mathbf{S}_{t}}\mathcal{I}(X_{i}:B_{t})$
@@
$\sum_{t=1}^{n}\sum_{X_{i}\in\mathbf{S}_{t}}S(X_{i}|B_{t})\geq$ $\frac{1}{m-1}\sum_{i<j}q^{X_{i,j}}(\rho_{A})+\frac{m}{2}S(A)$ (26)
- $\sum_{t=1}^{n}\sum_{X_{i}\in\mathbf{S}_{t}}\mathcal{I}(X_{i}:B_{t})-\frac{1}{m-1}\sum_{i<j}^{m}\tilde{f}^{X_{i,j}}.$
@@
$S(X_{i}|B_{t})+S(X_{j}|B_{s})\geq\left\{\begin{aligned} &q^{X_{i,j}}(\rho_{A}),\qquad t\neq s\\ &q^{X_{i,j}}(\rho_{A})+S(A|B_{t})-f^{X_{i,j}},t=s.\end{aligned}\right$.
@@
$\sum_{t=1}^{n}\sum_{X_{i}\in\mathbf{S}_{t}}S(X_{i}|B_{t})\geq$ $\frac{1}{m-1}\sum_{t=1}^{n}\frac{m_{t}(m_{t}-1)}{2}S(A|B_{t})$ (27)
$+$ $\frac{1}{m-1}\sum_{i<j}q^{X_{i,j}}(\rho_{A})$
- $\frac{1}{m-1}\sum_{t=1}^{n}\mathop{\sum}_{\begin{subarray}{c}X_{i^{\prime}},X_{j^{\prime}}\in S_{t}\\ i^{\prime}<j^{\prime}\end{subarray}}f^{X_{i^{\prime},j^{\prime}}},$
@@
$\sum_{i=1}^{m}S\left(X_{i}\mid B\right)\geq$ $\frac{1}{m-1}\sum_{i<j}q^{X_{i,j}}(\rho_{A})+\frac{m}{2}S(A|B)$
- $\frac{1}{m-1}\sum_{i<j}f^{X_{i,j}}+\max\left\{0,\kappa_{m1}\right\},$
where
$\kappa_{m1}=\displaystyle\frac{m}{2}\mathcal{I}(A:B)-\sum_{i=1}^{m}\mathcal{I}(X_{i}:B)+\frac{1}{m-1}\sum_{i<j}^{m}\left[f^{X_{i,j}}-\tilde{f}^{X_{i,j}}\right]$
If the cardinality of $\mathbf{S}_{t}$ is set to be 11 ($m_{t}=1$, $n=m$), Theorem 4 further simplifies to
@@ IV.1 Uncertainty relation for quantum coherence
Coherence is a fundamental issue in quantum physics and a crucial resource for quantum information processing. In 2014 Baumgratz et al. Baumgratz et al. (2014) introduced a framework for quantifying coherence. For a fixed orthonormal basis $\{|i\rangle\}^{d}_{i=1}$ of the $d$-dimensional Hilbert space $\mathcal{H}$, density matrices that are diagonal in this basis are termed incoherent. We denote the set of incoherent quantum states as $\mathcal{I}=\{\sigma|\sigma=\sum\limits_{i=1}^{d}\lambda_{i}|i\rangle\langle i|\}$.
For a bipartite system $AB$, an $A$-incoherent state (also known as an incoherent-quantum state Chitambar et al. (2016)) was proposed in Ma et al. (2016) with respect to the basis ($|i\rangle_{A}$}) of subsystem $A$ . This is also referred to as unilateral coherence Ma et al. (2017); Dolatkhah et al. (2019). We denote the set of $A$-incoherent states as $\mathcal{I}_{B|A}$,
@@
$C^{B|A}_{r}(\rho_{AB})=$ $\mathop{\mathrm{min}}\limits_{\sigma_{AB}\in\mathcal{I}_{B|A}}S(\rho_{AB}\|\sigma_{AB})=S(\rho_{AB}\|\Delta_{A}(\rho_{AB}))$
$=$ $S(\Delta_{A}(\rho_{AB}))-S(\rho_{AB}),$
@@
$\sum_{t=1}^{n}\sum_{M_{i}\in\mathbf{S}_{t}}C^{M_{i}}_{r}(\rho_{AB_{t}})\geq$ $\frac{1}{m-1}\sum_{i<j}Q^{M_{i,j}}$ (30)
$+$ $\sum_{t=1}^{n}\frac{m_{t}(m_{t}-2m+1)}{2(m-1)}S(A|B_{t})$
$+$ $\max\{0,\delta_{mn},\delta_{mn}^{{}^{\prime\prime}}\}.$
We employ an analogous approach to generate a random two-qutrit state $\rho_{AB}$. Unlike the three-qutrit case, $\rho_{AB}$ has nine eigenvalues and eigenvectors here. The procedure begins with generating a random $9\times 9$ real matrix $R_{1}$, from which we construct the Hermitian matrix $\mathcal{M}_{1}=D_{1}+(U_{1}^{T}+U_{1})+\mathrm{i}(L_{1}^{T}+L_{1})$. Here $D_{1}$ contains the diagonal elements of $\mathcal{R}_{1}$, while $U_{1}$ and $L_{1}$ represent its strictly upper and lower triangular components, respectively. Diagonalizing $\mathcal{M}_{1}$ yields nine normalized eigenvectors, which are then used to form a random unitary matrix $\mathcal{V}_{1}$. The remaining steps mirror the three-qutrit construction method, ultimately producing the desired two-qutrit state.
Example 3 Consider $10^{4}$ two-qutrit random states $\rho_{AB}$ and three measurements $M_{1}$, $M_{2}$ and $M_{3}$ with the corresponding eigenvectors $|\varphi^{1}_{1}\rangle=(\frac{\sqrt{2}}{2},\frac{\sqrt{2}}{2},0)^{T}$, $|\varphi^{1}_{2}\rangle=(-\frac{\sqrt{2}}{2},\frac{\sqrt{2}}{2},0)^{T}$, $|\varphi^{1}_{3}\rangle=(0,0,1)^{T}$, $|\varphi^{2}_{1}\rangle=(\frac{\sqrt{6}}{6},-\frac{\sqrt{6}}{6},\frac{\sqrt{6}}{3})^{T}$, $|\varphi^{2}_{2}\rangle=(-\frac{\sqrt{3}}{3},\frac{\sqrt{3}}{3},\frac{\sqrt{3}}{3})^{T}$, $|\varphi^{2}_{3}\rangle=(\frac{\sqrt{2}}{2},\frac{\sqrt{2}}{2},0)^{T}$, $|\varphi^{3}_{1}\rangle=(\frac{\sqrt{2}}{2},0,\frac{\sqrt{2}}{2})^{T}$, $|\varphi^{3}_{2}\rangle=(0,1,0)^{T}$, $|\varphi^{3}_{3}\rangle=(-\frac{\sqrt{2}}{2},0,\frac{\sqrt{2}}{2})^{T}$. In this scenario ($m_{t}=m=3$, $n=1$), our inequality (30) gives rise to
$C^{M_{1}}_{r}(\rho_{AB})+C^{M_{2}}_{r}(\rho_{AB})+C^{M_{3}}_{r}(\rho_{AB})$ $\geq\frac{3}{2}-\frac{3}{2}S(A|B)$ (31)
$+\max\{0,\delta_{31},\delta_{31}^{{}^{\prime\prime}}\},$
@@
$\delta_{31}^{{}^{\prime\prime}}=$ $-\frac{1}{3}\omega\beta+\frac{1}{2}S(A)+\frac{3}{2}\mathcal{I}(A:B)-\frac{1}{2}\sum_{i<j}Q^{M_{i,j}}$
$-\sum_{i}^{3}\mathcal{I}(M_{i}:B).$
Figure 4: Uncertainty relation for quantum coherence under three multiple measurements for $10^{4}$ randomly generated two-qutrit states. The horizontal axis represents the lower bound in inequality (31), while the vertical axis represents the value of $C^{M_{1}}_{r}(\rho_{AB})+C^{M_{2}}_{r}(\rho_{AB})+C^{M_{3}}_{r}(\rho_{AB})$. The red solid line is the proportional function with a slope of unity, displaying that the lower bound in inequality (31) is equal to $C^{M_{1}}_{r}(\rho_{AB})+C^{M_{2}}_{r}(\rho_{AB})+C^{M_{3}}_{r}(\rho_{AB})$. The green circles denotes the value of $C^{M_{1}}_{r}(\rho_{AB})+C^{M_{2}}_{r}(\rho_{AB})+C^{M_{3}}_{r}(\rho_{AB})$.
In Figure 4, the coherence-based uncertainty relation (31) is illustrated for $10^{4}$ randomly generated two-qutrit states and three measurements.